OpenStudy (calculusxy):

How would I graph 2x + 3y = 6

OpenStudy (calculusxy):

@jim_thompson5910

OpenStudy (calculusxy):

@satellite73

jimthompson5910 (jim_thompson5910):

if x = 0, then what is the value of y?

OpenStudy (calculusxy):

0 + 3y = 6 y = 2

jimthompson5910 (jim_thompson5910):

so (0,2) is one point on the graph

jimthompson5910 (jim_thompson5910):

if y = 0, then x = ??

OpenStudy (calculusxy):

x = 3

jimthompson5910 (jim_thompson5910):

(3,0) is the second point you need just 2 points to graph a linear equation

OpenStudy (calculusxy):

I mean I need help using it "algebraically"

jimthompson5910 (jim_thompson5910):

you mean solving 2x + 3y = 6 for y?

OpenStudy (calculusxy):

2x + 3y = 6 2x - 2x + 3y = 6 - 2x 3y = 6 - 2x

OpenStudy (calculusxy):

But I am having confusions solving it from there.

jimthompson5910 (jim_thompson5910):

to isolate y in 3y = 6-2x, divide both sides by 3

OpenStudy (calculusxy):

So it would be y = 2 - 2/3x ?

jimthompson5910 (jim_thompson5910):

yes \[\Large y = -\frac{2}{3}x + 2\]

OpenStudy (calculusxy):

Why did you switch it ?

OpenStudy (calculusxy):

Like is it necessary to do that?

jimthompson5910 (jim_thompson5910):

I placed it into y = mx+b form

jimthompson5910 (jim_thompson5910):

no it's not mandatory

OpenStudy (calculusxy):

Okay thanks :)

jimthompson5910 (jim_thompson5910):

np